# Understanding Nonlinear Collaboration between Human and AI Agents: A Co-design Framework for Creative Design

### 14 Jan 2024 v1

@@ 3.2. Experiment Design: A Formative Study
To gain insights from experts’ practice, we conducted a human-human collaborative design study. The formative study was comprised of three steps, introduction, modification, and interview. All sessions were recorded for subsequent analysis with the participants’ consent.
@@ 3.2.1. Participants
We recruited participants with a minimum of one year of design experience through a recruitment message posted on various online social media platforms. These individuals were further selected based on the quality of their design work. In total, our study involved 12 participants (7 women and 5 men), all of whom had their design artworks showcased in prominent events such as public activities, design competitions, or art exhibitions with a substantial number of attendees exceeding hundreds. While recruiting participants, we collected 5 graphic designs and the corresponding modification requests from candidates. These designs involved the topics of medicine, biology, literature, product design, game design, and so on. To ensure design diversity, we selected 3 out of 5 designs that exhibit variations in layout, target audience, and modification requests as formal study material.
@@ 3.2.2. Procedure
The study started with an overview of the collaborative modification process for single-page layouts, aiming to familiarize participants with the goal of this formative study. We explained the modification process as a collective effort, where individuals collaboratively devise concrete modification strategies based on initially vague requirements. To maintain coherence, the modification strategies needed to fall within the scope of single-page layouts.
In the modification session, two participants took part in the study simultaneously, with one assuming the role of the designer and the other acting as the design assistant (role-playing). The design assistant was asked to modify the designs submitted by the designer. After that, they switched roles and repeated the modification session which lasted around 25 minutes. The modification session was detailed as follows:
  • Content Understanding. The designer explain the content of the designs and demonstrat the associated requirements within 5 minutes. During that, the design assistant have the opportunity to interrupt and ask questions regarding the design content.
  • Independent Modification Strategies Brainstorming. After understanding the modification task, the design assistant is allotted 5 minutes to brainstorm potential modification strategies in the absence of the designer. The design assistant is permitted to sketch or write on Apple Freeform as part of the brainstorming process.
  • Synchronization. We then give the designers and design assistant 10 minutes to reach an agreement on the modification requirement and concrete strategies for each design.
  • Modificaiton Summarizing and Rating. After that, the design assistant is asked to provide an oral summarization of the optimization methods employed and describe the process of how they reached a consensus. After the summarization, the designer provide feedback on the design assistant using a 5-point rating scale, assessing the effectiveness of requirement analysis and fulfillment.
We then proceeded with a semi-structured interview with each pair of participants to gain deeper insights into the scenario and further explore users’ engagement in the future utilization of AI agents in design modification tools. We asked each participant the following questions. (i) What are the challenges you encountered when you collaborated with others in design presentations and during this experiment? What is your role during that process? (ii) What are the advantages and disadvantages of collaborative design modification compared to independent design modification? (iii) What assistance would you expect from AI agents during the process of modifying the design? Are there any tasks or aspects that you prefer to handle independently, considering that they might exceed the capabilities of AI assistants? If so, what are the reasons behind your preference? Also, several additional questions were tailored based on the participant’s performance during the modification session.
@@ 3.3. Expert Feedback and Discussion
In this session, we summarize experts’ feedback from the collected RA data in modification sessions and the interview recordings and propose five design requirements for the Human-AI Co-design framework.
@@
We employed two approaches to summarize the characteristics of requirements. Firstly, we analyzed the evolution of design requirements during the Modification Session. Secondly, we gathered feedback related to requirements analysis through Semi-structured Interviews. The process of requirement evolution in co-design is chaotic. When modifying designs, requirements evolve, and users gain more insight into what they need in the process of continuous trial and error. The inherent nonlinearity of requirements arises brought about three challenges: requirement ambiguity, requirement changes, and requirement conflicts. Participants who played the role of designers in the first session were labeled as A, whereas the partner participants were labeled as B. Specifically, $A_{ij}$ refers to the designers in the $i$-th group with regards to the $j$-th design. For further details, the relevant words of experts are provided in Appendix A.
@@
  • Requirement Ambiguousness. The initial version of design requirements was often underdeveloped and prone to ambiguity, which poses the most significant challenge. In $A_{12}$ and $B_{23}$, designers voiced dissatisfaction without providing any optimization intentions. In other cases, optimization intentions were expressed in a general manner. $A_{32}$ and $A_{33}$ described the visual effect of a picture as a vague sensation rather than a visible outcome. “Identifying” ($A_{11},A_{23},A_{41},B_{61},B_{63}$) and “improving” ($A_{42},A_{53},A_{63},B_{12},B_{31},B_{42}$) pages are among the most commonly encountered ambiguous descriptions. Compounding the issue, some participants ($A_{11},A_{23},B_{61},B_{12},B_{41},B_{42}$) failed to clarify the specific parts of the designs to be optimized.
@@
  • Requirement Changes. While a small number of the requirements ($B_{32}$, $B_{52}$) were satisfied at once, the majority of the designs (34/36) were iterated in a trial-and-error manner. Some shifts in design requirements might be generated when correcting inaccurate descriptions. Some requirements undergo iterations, while new requirements emerge in the designers’ minds, as a result of interpersonal dynamics within the design team. In the course of communication, members of the design group make compromises to harmonize their perspectives and achieve a collective agreement on design requirements.
@@
  • Requirement Conflict. Requirements were constantly subject to negotiation and could often clash with one another. We identified three types of conflicts: (i) Conflicts between asynchronous requirements, which arise naturally as designers change their minds, normally designers focus on previous ones; (ii) Conflicts between individuals, which happen every now and then in collaboration, as different people see things from different perspectives; and (iii) Conflicts within the design itself, which may involve factors such as functionality, aesthetics, budget, time constraints, user preferences, technical feasibility, or ethical considerations. Finding a balance between conflicting requirements involves exploring alternative design options, presenting selective outcomes for feedback, and iterating on the design based on negotiation. It is crucial to prioritize the core requirement while accommodating as many perspectives as possible.
@@
The co-design process is an imaginative voyage of discovery. According to the interview, the presence of another designer brings forth a multitude of distinctions and divergences in many aspects. The pros (+) and cons (-) of the nonlinear human-human co-design process are as follows:
    Inspiration to design in alternative ways. In practice, every designer’s comprehension of design is inherently restricted. Despite their extensive knowledge and skills, designers’ approaches to modification are still confined by their understanding of the requirements at hand. The inclusion of an extra viewpoint from a collaborator can stimulate creativity. Our study has demonstrated that unanticipated inspiration from others can enable the achievement of design goals through alternative methods.
    Collaboration accelerates the design progress under timely communication. When designers work on refining their designs independently, they may become preoccupied with minor details. In contrast, when collaborating, the need for prompt communication necessitates swift responses. Consequently, ideas are rapidly exchanged and iterated upon until both designers and design assistants reach a mutual agreement.
    Misunderstandings are exposed and fixed swiftly through Q&A. Frequent sync allows people to detect and react to misunderstandings timely. During communication with their design assistants, designers are actively contemplating how to convey the primary concept of their designs. By observing their partners’ reactions, they become aware of potential misunderstandings regarding the designs. In turn, design assistants gain a proper understanding of the designs as any misconceptions are rectified during the conversation.
    Real-time outcomes lacking in human-human collaboration. We also notice the lack of real-time outcomes in human-human collaboration. Participants could only imagine a visual outcome according to the texts rather than finish these modifications in a short time.
Based on the above discussion, we propose the following design requirements (table. 1) for a human-centered nonlinear framework to support better human-AI co-design.
@@
Table 1. Design requirements for Human-AI Co-design systems.
@@ 4. A Human-AI Co-design Framework for Creative Design
To support the outlined design requirements, we formulate a nonlinear human-AI co-design framework (Fig.4) that incorporates specific actions for AI agent. The framework is implemented in a proof-of-concept (POC) prototype called OptiMuse, where we focus on user interface (UI) designs. Two primary components of the nonlinear framework will be detailed in the subsequent sections.
4.1. AI Actions in the Nonlinear Co-design Framework
@@
Drawing from our formative study (Sec. 3), we recognize the various expectations designers hold for their collaborators. Realizing these expectations can significantly enhance the nonlinear co-design experience, a facet not yet fully actualized in current human-AI collaborations (lubart2005can, ; wang_literature_2017, ; muller-wienbergen_leaving_2011, ; negrete2014apprentice, ; kantosalo2016modes, ; guzdial2019interaction, ; rezwana2022understanding, ). These expectations are primarily centered around five areas, corresponding to DR 1-5, namely: flexible communication, clarification, alternative solutions, operational strategies, and visual outcomes. Concerning the first two design requirements, designers are more inclined to confidently express ambiguous descriptions when they have opportunities to provide explanations and rectify any misunderstandings. To facilitate and promote the use of ambiguous commands, a communication procedure is essential (shi2023Understanding, ). To fulfill user expectations, we emphasize the necessity of the following AI agent actions within the communication process:
  • A1: Guiding and supporting users in refining ambiguous descriptions. (Corresponding to DR1, DR2)
@@
  • A2: Providing reminders to users when requirements change or conflicts arise. (DR1)
  • A3: Demonstrating alternative explanations and descriptions of users’ commands. (DR2)
@@
Regarding the third design requirement, designers present a range of design alternatives to meet requests, drawing from their expertise or existing samples. They derive inspiration through trial and error as they navigate among these design choices. Various options were exemplified through a combination of textual descriptions, verbal explanations, sketches, supported gestures, and comprehensive visual results (shi2023Understanding, ). Navigating, summarizing, and discarding among these various options are central to the nonlinear co-design process (Inakage2007collective, ; tan2022conflict, ; Cheliotis2014remix, ; Oehlberg2015remix, ; yue2019improving, ). The following actions were frequently observed and, as a result, documented as required actions of AI agents for multiple selections corresponding to DR3:
  • A4: Revising users’ requirements and offering different or additional choices. (DR3)
  • A5: Analyzing choices to detect distinctions and select superior ones. (DR3)
  • A6: Modifying the optimal choice for enhanced results. (DR3)
  • A7: Making revisions to multiple options and conducting subsequent comparisons. (DR3)
  • A8: Combining elements from diverse options to enhance results. (DR3)
@@
In alignment with the design requirements addressing the cons of human-human collaboration, it is noted that:
  • A9: AI agent outputs should be operational strategies. (DR4)
  • A10: AI agent can preview the visual effect of operational strategies. (DR5)
@@
These then proposed AI actions are specifically crafted to address Design Requirements 1-5, striving to provide designers with a nonlinear experience that more closely resembles human-human collaboration process.
Figure 3. To achieve a final design outcome, most AI agents execute a series of gradual and precise commands.
@@
Figure 4. We propose a nonlinear human-AI co-design framework characterized by facilitating a communication process before producing visual results and accommodating versatile actions for multiple choices.
Figure 5. OptiMuse overview: (a) The preview shows current outcomes; (b) The primary communication process in OptiMuse occurs through dialogues; (c) OptiMuse generates three distinctive choices in the form of comprehensive visual outcomes each time.
4.2. User Interface of OptiMuse: a Proof-of-concept Prototype
To facilitate effective interaction (Fig. 6) between designers and AI, we have developed the user interface of OptiMuse (Fig. 5). This UI comprises three primary components: a preview window, a dialogue box, and selective choices. The preview window displays the current output of OptiMuse, offering users a dynamic visual representation (A10) of the design as it evolves through interaction. Regarding the dialogue component, this is where designers communicate with the AI agent. Study I identified three dominant forms of flexible communication: text, speech, and draft. However, due to the higher complexity in speech and draft interactions, text-based communication is more common in popular AI-driven tools (copilot, ; midjourney, ). Designers are accustomed to engaging in prompt-based communication with AI agents and expecting visual outputs due to the prevalence of TTIs (e.g. Imagen (google2022imagen, ), Parti (google2022parti, ), and DALL-E 2 (openai2022dalle2, ; openai2022dalleblog, )). Consequently, we opted for text-based dialogue as the basic form of communication process.
@@ 4.2. User Interface of OptiMuse: a Proof-of-concept Prototype
Apart from the required AI agents actions (A1-3), the dialogue regulations are as follows (see Appendix B for more details) : designers engage in conversations with OptiMuse to align on optional strategies before they generate corresponding visual outcomes. They are mandated to transmit the signal “ Let’s Have a try! ” whenever they wish, indicating to OptiMuse to generate visual outcomes. Subsequently, they can issue commands and make reference to one or multiple options provided by it to continue their interactions based on the available choices. If they refer to only one option, the current outcome will be updated to reflect their preference. In addition, if they are dissatisfied with any of the provided choices, they have the option to signal “ NO ” along with a new command to continue the discussion based on the existing outcome. Once they are content with the current outcome, they are required to transmit “ THX OptiMuse! ” to signify the conclusion of the design process.
@@
As for the selective choices, OptiMuse’s UI presents the various visual outcomes provided by the AI agent (A10). This design addresses the limitations of traditional human-human collaboration (A9,10) while preserving the nonlinear nature of the process. The selections encompass different design parameter options, interpretations of requirements, modification strategies, and combinations thereof (A4). To balance attention and choice overload (chernev2015choice, ), we limit to three selective options to highlight the features of multiple choices. OptiMuse generates three distinct visual outcomes for each selection, with additional choices potentially emerging during dialogue. Users can request textual descriptions of these choices for review. The content of the selective choices is determined by the current outcome, the dialogue that occurs after the last “ Let’s have a try! ”, and OptiMuse’s design knowledge. To foster creativity, OptiMuse not only executes users’ commands but also has the capacity to modify other aspects of the design, as long as these modifications align with design principles and do not contradict users’ requirements. If required by users, OptiMuse can execute A5-8 on the selective choices.

### 20 Mar 2024 v3

@@ 3.2. Experiment Design: A Formative Study
To gain insights from experts’ practice, we conducted a human-human collaborative design study that was recorded for subsequent analysis with participants’ consent.
@@ 3.2.1. Participants
We recruited participants with a minimum of one year of design experience through a recruitment message posted on various online social media platforms. These individuals were further selected based on the quality of their design work. In total, our study involved 12 participants (7 women and 5 men). All of them had their designs showcased in prominent events such as public activities, design competitions, or art exhibitions with a substantial number of attendees exceeding hundreds. When recruiting participants, we collected 5 graphic designs and their modification requests from candidates. These designs involved topics of medicine, biology, literature, product design, game design, and so on. To ensure design diversity, we selected 3 out of 5 designs that exhibit variations in layout, target audience, and modification requests as formal study material.
@@ 3.2.2. Procedure
The study started with an overview of the collaborative modification process. We explained the modification process as a collective effort, where individuals collaboratively devise concrete modification strategies based on initially vague requirements. To maintain coherence, the modification strategies needed to fall within the scope of single-page layouts.
In the modification session, two participants took part in the study simultaneously, one playing the role of designer and the other of design assistant. The design assistant was asked to modify the designs submitted by the designer. After that, they switched roles and repeated the modification session. The 25-minute modification session was detailed as follows:
  • Content Understanding. The designer explains the content of the designs and demonstrates the requirements within 5 minutes while the design assistant has the opportunity to interrupt and ask questions about design content.
  • Independent Modification Strategies Brainstorming. After understanding the modification task, the design assistant has 5 minutes to brainstorm potential modification strategies independently. The design assistant is allowed to sketch or write on Apple Freeform as part of the brainstorming process.
  • Synchronization. Then the designers and design assistant have 10 minutes to reach a consensus on the modification requirement and specific optimization strategies for each design.
  • Modificaiton Summarizing and Rating. The design assistant is asked to provide an oral summarization of the optimization strategies employed and describe how they reach a consensus. After that, the designer provides feedback on the design assistant using a 5-point rating scale, assessing the effectiveness of requirement analysis and fulfillment.
We then proceeded with a semi-structured interview to gain deeper insight into the scenario and further explore users’ engagement in the future utilization of AI agents in design modification tools. We asked each participant the following questions: (1) What are the challenges you encountered when collaborating with others in design modification and during this experiment? What is your role during that process? (2) What are the advantages and disadvantages of collaborative design compared to independent design? (3) What assistance do you expect from AI when modifying the design? Are there any tasks or aspects that you prefer to handle independently? If so, what are the reasons behind your preference? Several additional questions were tailored based on the participant’s performance during the modification session.
@@ 3.3. Expert Feedback and Discussion
Based on collected RA data in modification sessions and interview recordings, we summarize experts’ feedback on the characteristics of requirements, as well as the pros and cons of the nonlinear human-human co-design process. Subsequently, we outline five design requirements (Table 1) for Human-AI Co-design.
@@
We employed two approaches to summarize the characteristics of requirements: (1) We analyzed the evolution of design requirements during the Modification Session; (2) We gathered feedback related to requirements analysis through Semi-structured Interviews. The process of requirement evolution is chaotic. When modifying designs, requirements evolve, and users gain more insight into what they need in trial and error. The inherent nonlinearity of requirements are reflected as: requirement ambiguity, requirement changes, and requirement conflicts. Participants who played the role of designers in the first session were labeled as A, whereas the partner participants were labeled as B. Specifically, $A_{ij}$ refers to the designers in the $i$-th group with regards to the $j$-th design. For further details, the relevant words of experts are provided in Appendix A.
@@
  • Requirement Ambiguity. The initial version of design requirements was often underdeveloped and prone to ambiguity. In $A_{12}$ and $B_{23}$, designers voiced dissatisfaction without providing any optimization intentions. In other cases, optimization intentions were expressed in a general manner. $A_{32}$ and $A_{33}$ described the visual effect of a picture as a vague sensation rather than a visible outcome. “Identifying” ($A_{11},A_{23},A_{41},B_{61},B_{63}$) and “improving” ($A_{42},A_{53},A_{63},B_{12},B_{31},B_{42}$) pages are the most frequent ambiguous descriptions. Compounding the issue, some participants ($A_{11},A_{23},B_{61},B_{12},B_{41},B_{42}$) failed to clarify specific parts of the designs to be optimized.
@@
  • Requirement Changes. While a few requirements ($B_{32}$, $B_{52}$) were fulfilled in a single iteration, the majority of designs (34 out of 36) underwent trial-and-error iterations. Some shifts in design requirements are generated when correcting inaccurate descriptions. Some requirements undergo iterations, while new requirements emerge in designers’ minds, as a result of interpersonal dynamics within the design team. In communication, cooperators make compromises to harmonize their perspectives and reach a consensus on design requirements.
@@
  • Requirement Conflict. Requirements are constantly subject to negotiation and often clash with one another. We identify three types of conflicts: (i) Conflicts between asynchronous requirements, which arise naturally as designers change their minds, normally designers focus on current ones; (ii) Conflicts between individuals, which frequently happen because different people see things from different perspectives; and (iii) Conflicts within the design itself, which involve factors such as functionality, aesthetics, budget, time constraints, user preferences, technical feasibility, and ethical considerations. Striking a balance between conflicting requirements involves exploring alternative design options, presenting selective outcomes for feedback, and iterating on the design based on negotiation. It is crucial to prioritize the core requirement while accommodating as many perspectives as possible.
@@
The co-design process is an imaginative voyage of discovery. According to the interview, the presence of another designer brings forth a multitude of distinctions and divergences in many aspects, which brings about both pros (+) and cons (-) of the nonlinear human-human co-design process.
    Inspire to design in alternative ways. In practice, every designer’s comprehension of design is restricted. Despite their extensive knowledge and skills, designers’ approaches to modification are confined by their understanding of the requirements. The inclusion of an extra viewpoint from a collaborator can stimulate creativity. Our study demonstrates that unanticipated inspiration from others enables achieving design goals through alternative methods.
    Accelerate design progress in timely communication. When designers refine designs independently, they sometimes are stuck in minor details. In contrast, when collaborating, the need for prompt communication necessitates swift responses. Consequently, ideas are rapidly iterated until both designers and design assistants reach a consensus.
    Expose and fix misunderstandings through Q&A. Frequent sync allows people to detect and react to misunderstandings timely. During communication with their design assistants, designers are actively thinking about how to convey the primary concept of designs. By observing their partners’ reactions, they become aware of potential misunderstandings on the designs. In turn, design assistants gain a proper understanding as misconceptions are rectified during the conversation.
    Lack real-time outcomes. Despite all the pros, designers report a lack of real-time outcomes in human-human collaboration. Due to the time-consuming manual modification, participants can only imagine a visual outcome according to words and sketches, which takes a toll on accuracy.
To strengthen the pros and address the cons, we further summarize five design requirements (table  1) for a human-centered nonlinear framework to support better human-AI co-design.
Table 1. Design requirements for Human-AI Co-design systems.
@@ 4. A Human-AI Co-design Framework for Creative Design
To fulfill the design requirements, we formulate a nonlinear human-AI co-design framework that incorporates specific actions for AI agents. The framework is implemented in a POC called OptiMuse, where we focus on user interface design. Two primary components of the nonlinear framework are detailed in the subsequent sections.
4.1. Expected Actions for AI Agents
@@
Drawing from the formative study (Sec. 3), we summarize 10 expected actions designers hold for collaborators. Realizing these expectations can significantly enhance the co-design experience, a facet not yet fully actualized in current human-AI collaborations (Lubart, 2005; Wang and Nickerson, 2017; Müller-Wienbergen et al., 2011; NEGRETE YANKELEVICH and MORALES ZARAGOZA, 2014; Kantosalo and Toivonen, 2016; Guzdial and Riedl, 2019; Rezwana and Maher, 2022). These expected actions, which are denoted as A1-A10, are primarily centered around DR 1-5: flexible communication, clarification, alternative solutions, operational strategies, and visual outcomes.
Concerning the first two design requirements, designers are more inclined to confidently express ambiguous descriptions if they have opportunities to provide explanations and rectify misunderstandings. To facilitate the use of ambiguous commands, a communication process before AI agents generate outcomes is essential (Shi et al., 2023). To be specific, we emphasize the necessity of the following AI agent actions within the communication process:
  • A1
    Guide and support users in refining ambiguous descriptions. (Corresponding to DR1, DR2)
  • A2
@@
    Provide reminders to users when requirements change or conflicts arise. (DR1)
  • A3
    Demonstrate alternative explanations and descriptions of users’ commands. (DR2)
@@
Regarding the third design requirement, design assistants present multiple selections to meet requests, based on design knowledge or existing examples. These selections were presented through a combination of textual descriptions, verbal explanations, sketches, supported gestures, and comprehensive visual results (Shi et al., 2023) as design choices. Designers derive inspiration when navigating among these design choices. Navigating, summarizing, and discarding among the design choices are central to the nonlinear process of collective creativity  (Inakage, 2007; Tan et al., 2022; Cheliotis et al., 2014; Oehlberg et al., 2015; Yue et al., 2019). The following actions were frequently observed in the formative study and, therefore, documented as required actions of AI agents in generating multiple selections:
  • A4
    Revise users’ requirements and offer different or additional choices. (DR3)
  • A5
    Review choices to identify difference or superior options. (DR3)
  • A6
    Modify the optimal choice for enhanced results. (DR3)
  • A7
    Make revisions to multiple options and conduct subsequent comparisons. (DR3)
  • A8
    Merge elements from different options. (DR3)
@@
In alignment with the fourth and fifth design requirements addressing the cons of human-human collaboration, two required actions of AI agents are notable for outputs:
  • A9
    Clarify operational strategies. (DR4)
  • A10
    Preview the visual effect of operational strategies. (DR5)
@@
These ten proposed AI actions are specifically crafted to address Design Requirements 1-5, striving to provide designers with a nonlinear experience that more closely resembles the flexibility and inspiration in the human-human collaboration process (Fig. 3).
@@
Figure 3. In the traditional linear human-AI collaboration, AI agents execute a series of gradual and precise commands to achieve final design outcomes. We propose a nonlinear human-AI co-design framework characterized by facilitating a communication process before producing visual results and accommodating versatile actions for multiple choices.
Figure 4. User interactions in OptiMuse follow these steps: (1) users review slides on the preview window; (2) users input commands and engage in rule-based conversation with OptiMuse; (3) OptiMuse generates alternatives as selective choices upon reaching convergence with users; (4) users remix choices through rule-based conversation; (5) users are satisfied with the choices and choose one option after several rounds of steps 1 to 4 (Notes: A1-10 refer to the AI agent actions defined in Sec. 4.1).
4.2. OptiMuse: a Proof-of-concept Prototype
To facilitate effective interaction (Fig. 4) between designers and AI, we have developed the user interface of OptiMuse (Fig. 7). This user interface comprises three primary components: a preview window, a dialogue box, and selective choices.
The preview window displays the latest output of OptiMuse, offering users a dynamic visual representation of the design (Corresponding to A10) as it evolves through interaction.
@@ 4.2. OptiMuse: a Proof-of-concept Prototype
Regarding the dialogue box, this is where designers communicate with the AI agent. Study I identified three dominant forms of flexible communication: text, speech, and draft. However, due to the higher complexity in speech and draft interactions, text-based communication is more common in popular AI-driven tools (Microsoft, 2023; Midjourney, 2022). Designers are accustomed to engaging in prompt-based communication with AI agents and expecting visual outputs due to the prevalence of prompt-based text-to-image models (e.g. Imagen (Google, 2022a), Parti (Google, 2022b), and DALL-E 2 (OpenAI, 2022a, b)). Consequently, we opted for text-based dialogue as the basic form of communication process. To allow the required actions of AI agents (A1-3) for the communication process, we designed specific dialogue regulations in OptiMuse (see Appendix B for detailed rules for AI agents) : designers engage in conversations with OptiMuse to align on optional strategies before they generate corresponding visual outcomes. They are mandated to transmit the signal “ Let’s Have a try! ” whenever they wish, indicating to OptiMuse to generate visual outcomes. Subsequently, they can issue commands and make reference to one or multiple options provided by it to continue their interactions based on the available choices. If they refer to only one option, the current outcome will be updated to reflect their preference. In addition, if they are dissatisfied with any of the provided choices, they have the option to signal “ NO ” along with a new command to continue the discussion based on the existing outcome. Once they are content with the current outcome, they are required to transmit “ THX OptiMuse! ” to signify the conclusion of the design process.
@@
As for the selective choices, OptiMuse presents the selectable visual outcomes (A9,10) provided by AI agents. This design preserves the nonlinear nature of the creative design process while addressing the limitations of traditional human-human collaboration. The selections encompass different design parameter options, interpretations of requirements, modification strategies, and combinations thereof (A4). To balance attention and choice overload (Chernev et al., 2015), we limit to three selective options to highlight the features of multiple choices. OptiMuse generates three distinct visual outcomes for each selection, with additional choices potentially emerging during dialogue. Users can request textual descriptions of these choices for review (A5). The content of the selective choices is determined by the current outcome, the dialogue that occurs after the last “ Let’s have a try! ”, and OptiMuse’s design knowledge. To foster creativity, OptiMuse not only executes users’ commands but also has the capacity to modify other aspects of the design, as long as these modifications align with design principles and do not contradict users’ requirements. If required by users, OptiMuse can perform further operations of reviewing (A5), modifying (A6), comparing (A7), and remixing (A8) the selective choices. Together, the user interface design of OptiMuse allows for natural, nonlinear interactions between human and AI agents (Fig. 4) to perform tasks in creative design.